did I do this right if not can you tell me what I did wrong
2x + y = 5
x − 2y = 10
y = 5 − 2x
x − 2(5 − 2x) = 10
x − 10 + 4x = 10
5x − 10 = 10
5x = 0
x = 0
2(0) + y = 5
y = 5

Answer by rothauserc(4718)

The next step after 5x-10=10 should be
5x=20 (add 10 to both sides of =)
and x=4
then (2*4) + y = 5 and y = -3